Given the balanced equation:
4CuO + CH4 -----> 4Cu + 2H2O + CO2

49 grams of cupric oxide was used to produce a 73% reaction yield. Determine the mass of carbon dioxide produced.
1
Expert's answer
2020-12-01T14:24:19-0500

"4CuO + CH4 ---> 4Cu + CO2 + 2H2O"

1) n(CuO)=m/M= 49/(64+16)=0,6125 mole

2) n(CO2)=n(CuO)/4=0,1531 mole

3) m(CO2)=n(CO2)*M(CO2)*reaction yield=0,1531*(12+32)*0,73=4,92g

Answer: 4,92 g